A ball is thrown horizontally from a cliff such that it strikes the ground after 5 seconds. The line of sight from the point of projection to the point of projection of hitting makes an angle of 37 degrees with the horizontal. What is the initial velocity of projection?
 

Grade:11

1 Answers

Arun
6 years ago
Dear Aditi
When you make the diagram you will see
Tan 37 = h/R
3/4 = h/R
R = 4 h/3
Also
T = \LARGE \sqrt(2h/g)
5 = \LARGE \sqrt(2h/g)
25 = 2h/g
Put g = 10
h = 125 metres
 
Now
R = u*T
u = R/T = 4*125/3*5
u = 100/3 m/sec
 
Regards
Arun (askIITians forum expert)